Calculate the radius and energy of the electron in the nth orbital in the hydrogen atom


1
Expert's answer
2021-03-08T08:21:57-0500

The radius and energy of the electron in the nth orbital in the hydrogen atom can be calculated in a semiclassical approximation using Bohr's postulates.

Start from the classical case. The electron is held in a circular orbit by electrostatic attraction. The centripetal force is equal to the Coulomb force:


"{m_\mathrm{e} v^2\over r} = {Zk_\mathrm{e} e^2 \over r^2}"

where "m_e" is the mass of the electron, "v" is its speed, "r" is the radius of the orbit, "Z" is the atom's atomic number ("Z = 1" for hydrogen), "k_e" is the Coulomb constant, and "e" is the charge of the electron.

Thus, the speed is:


"v = \sqrt{ k_\mathrm{e} e^2 \over m_\mathrm{e} r}"

and the energy is:


"E= -{1\over 2} m_\mathrm{e} v^2"

Now, let's use the Bohr's assumption - angular momentum is an integer multiple of "\hbar" :


"m_evr = n\hbar"

Substituting the expression for the velocity into the last expression gives an equation for r in terms of n:


"m_{\text{e}}\sqrt{\dfrac{k_{\text{e}}e^2}{m_{\text{e}}r}}r=n\hbar\\
r_n = {n^2\hbar^2\over k_\mathrm{e} e^2 m_\mathrm{e}}"

The same procedure for the energy gives:


"E_n = -{k_\mathrm{e} e^2 \over 2r_n } = - { k_\mathrm{e}^2 e^4 m_\mathrm{e} \over 2\hbar^2 n^2}"

Answer. "E_n= - { k_\mathrm{e}^2 e^4 m_\mathrm{e} \over 2\hbar^2 n^2}, r_n = {n^2\hbar^2\over k_\mathrm{e} e^2 m_\mathrm{e}}".